OpenStudy (anonymous):

The energy of a particular color of green light is 3.70E-22 kJ. The wavelength of this light is nanometers. (109 nm = 1 m)

OpenStudy (anonymous):

so i first converted kj to k...3.70e-22kj*(1/1j)=3.70e-22k.....then.....3.70e-22k*(1000m/1k)=3.7e-19m...then....3.7e-19m*(10^9nm/1m)=3.7e-10

OpenStudy (anonymous):

what did i do wrong???

OpenStudy (accessdenied):

I'm not sure I follow the conversions you are trying to do here. kJ is kiloJoules, where kilo is just a prefix. Converting from kJ to k does not seem to make sense. We ought to convert from kJ to J and then apply a formula between Eenrgy and wavelength to solve for wavelength. Then convert wavelength from meters to nanometers...

OpenStudy (anonymous):

oh so i have to use E=hv

OpenStudy (accessdenied):

Yes. Although v is frequency I believe, so we need just one more: c = (Lambda) v (c the speed of light, Lambda is wavelength, and v is frequency. This lets us convert between frequency and wavelength. Or the cumulative formula is just E = (h c)/ (Lambda)

OpenStudy (accessdenied):

\( c = \lambda v \) \(E = h v = \dfrac{h c}{\lambda} \)

OpenStudy (anonymous):

okay hold on im going to work it out

OpenStudy (anonymous):

so i converted kj to j and i got 3.7e-19j would this be v or λ, we know that c=2.9979x10^8m/c

OpenStudy (accessdenied):

that would be energy: "The energy of a particular color of green light is..." So that goes in for E. We want to find Wavelength, which is lambda. One way is by finding frequency v and then using the c = lambda * v formula to switch. Otherwise, we can just use this formula and find lambda directly.: E = hc / (lambda)

OpenStudy (anonymous):

but we dont have λ

OpenStudy (accessdenied):

Lambda is what we want to find, yes? That is wavelength, which the problem is asking for. h and c are just constants. And we have E given in the problem. Lambda is the only variable in the equation, so we can solve for it. :)

OpenStudy (anonymous):

okay v=5.36e-7?

OpenStudy (anonymous):

λ=5.59e14

OpenStudy (anonymous):

is this right

OpenStudy (accessdenied):

Hm.. 5.36 * 10^(-7) looks like the value for lambda I got. \( E = \dfrac{hc}{\lambda} \) \( 3.7 10^{-19} = \dfrac{(6.626*10^{-34}) (2.9979*10^8)}{\lambda} \)

OpenStudy (accessdenied):

3.7 * 10^(-19) , not 3.710.

OpenStudy (anonymous):

wait i used \[\frac{ hc }{ v }\]

OpenStudy (accessdenied):

If you solve here: \( 3.7 * 10^{-19} = (6.626 * 10^{-34} ) \nu \) \( \nu = \dfrac{3.7 *10^{-19}}{6.626 * 10^{-34}} \) comes out to 5.58 * 10^(14). I think you just reversed their roles.

OpenStudy (anonymous):

thats how i got v=5.36e-7

OpenStudy (accessdenied):

Yes; you effectively already had lambda, just switch v for lambda really. Then that would be wavelength in meters. Convert it to nanometers for the final answer. :)

OpenStudy (anonymous):

so you did E=hv....so when they give you E=hv=hc/v (they are 2 differeny equations not have to be used together?

OpenStudy (accessdenied):

Hmm... The original equation that I am looking at is this: \( E = h \nu \) \(E\) is energy in Joules, \(h\) is Planck constant in Joule-seconds. \( \nu \) is frequency in cycles per second. \( c = \lambda \nu \) \( \nu = \dfrac{c}{\lambda} \) This separate equation relates wavelength \(\lambda\) (meters) and frequency \(\nu \) to the speed of light. Substitution \( \nu = \dfrac{c}{\lambda} \) into the first equation, we have: \( E = h \dfrac{c}{\lambda} \) The usefulness of either equation is based on whether you have/need wavelength or frequency. In this case, we need wavelength, so we use the second equation. What I said previously is that technically, you *can* use the first one and find frequency, but you need to then convert to wavelength using the middle equation.

OpenStudy (anonymous):

oh okay that makes much more sense...wow thank you so much!

OpenStudy (accessdenied):

Glad to help! :) Were you able to find the answer? I am coming up with about 530 nm, which checks out according to the visible light spectrum.

OpenStudy (anonymous):

im doing it hold on

OpenStudy (anonymous):

λ=5.37e-7....i got 537

OpenStudy (accessdenied):

Yep, sounds good to me. :)

OpenStudy (anonymous):

its rights!!! yess thank you... could i ask you another question in a couple of mins?

OpenStudy (accessdenied):

I may have to sleep very soon (it's late here, 12am). I'll see what I can do. :)

OpenStudy (anonymous):

thank you...let me know then i could ask you tom

OpenStudy (accessdenied):

Is it a similar question to this one?

OpenStudy (anonymous):

yup

OpenStudy (accessdenied):

Should be fine now then? That shouldn't take too long.

OpenStudy (anonymous):

okay ill let you

OpenStudy (anonymous):

what are the units for frequency?

OpenStudy (anonymous):

A local FM radio station broadcasts at an energy of 6.52E-29 kJ/photon. Calculate the frequency at which it is broadcasting. Frequency = MHz (1 MHz = 106 sec -1) i got 3.03e-6 MHz

OpenStudy (accessdenied):

What was your process to obtain that answer?

OpenStudy (anonymous):

i converted kj to j.....\[\frac{ 6.52e-29kj }{photon }*\frac{ 1000j }{ ?1kj }=6.52e-26 j/photon\] then \[6.52e-26=\frac{ 6.626e-34*2.9979e8 }{ v }\] then \[v=\frac{ 1.98e-25 }{ 6.52e-26 }\] ....v=3.036809816

OpenStudy (anonymous):

with 3.036809816/10^6=3.03e-6MHz

OpenStudy (accessdenied):

I believe this is the same error from earlier... in that equation, v is in place of where wavelength is. \( E = \dfrac{h c}{\color{blue}\lambda} \) This is the equation with wavelength. \( E = h \color{red}{\nu} \) This is the equation with frequency.

OpenStudy (anonymous):

okay i got 98

OpenStudy (accessdenied):

98 MHz looks good to me now. It seems once you have the correct equation, you do fine with the math work. :)

OpenStudy (anonymous):

thank you so much for explaining it to me i was so confused before...

OpenStudy (accessdenied):

Yep, glad to help again. :D I will be heading off to bed now if that is all. Best of luck and gnight!

OpenStudy (anonymous):

good night...hope to talk to you again!